# Dynamical symmetry breaking in Georgi-Glashow chiral-gauge theories

### v1

@@ IV Patterns of symmetry breaking
$\partial_{t}\bar{\lambda}_{i}\propto$  $\,2\,\bar{\lambda}_{i}+\boldsymbol{c}_{{\rm A},i}\cdot\alpha_{g}^{2}+\boldsymbol{c}_{{\rm B},ij}\cdot\alpha_{g}\,\bar{\lambda}_{j}$
$\hskip 85.35826pt+\boldsymbol{c}_{{\rm C},ijk}\cdot\bar{\lambda}_{j}\bar{\lambda}_{k}+\ldots\,,$ (8)
where $\alpha_{g}=g^{2}/4\pi$ and we have only explicitly shown the lowest-order contributions. As the theory flows away from the UV Gaussian fixed point (the free theory), quantum fluctuations generate all kind of operators allowed by the symmetries via gauge interactions. More precisely, the four-fermion operators are triggered through gauge-boson-mediated box diagrams encoded in $\boldsymbol{c}_{{\rm A},i}$ in (8). Once generated, these feed into their own flows via the terms proportional to $\boldsymbol{c}_{{\rm B},ij}$, and the $\boldsymbol{c}_{{\rm C},ijk}$ terms control the self-interaction structure. The indices $j,k$ here, range over all channels including $i$ that corresponds to the basis in (6). Since fermionic self-interactions are triggered by gauge fluctuations, they may be interpreted as higher-order gauge corrections, with the various dressings corresponding to distinct tensor channels. For further details, see Braun (2012); Goertz et al. (2024); Dupuis et al. (2021) and references therein. The coefficients encode the color and global symmetries of the theory and display a non-trivial dependence on $N_{c}$ and $N_{\textrm{gen}}$, resulting from tracing over the internal structure of the theory. Finally, the first term in (8) accounts for the canonical dimension of the operator.
In total, the self-interaction terms $\boldsymbol{c}_{{\rm C},iii}$ yield a parabola shape (see Figure 4 and the respective particular example in Appendix B) of the flow as a function of its respective dressing which, depending on the sign of the coefficient can yield an upright or inverted shape. Moreover, for small $\alpha_{g}$ the flow of the dressings displays an IR attractive fixed point which in the classical limit ($k\to\infty$) reduces to the Gaussian one, $\alpha_{g}^{*}=0$ and $\bar{\lambda}_{i}^{*}=0$.
@@
As the gauge dynamics evolve towards the IR, the parabola shape is shifted upwards or downwards depending on the sign of the $\boldsymbol{c}_{{\rm A},i}$ coefficients, which read
[…]
Note that none of these depend on $N_{\textrm{gen}}$, but only on $N_{c}$ (see Figure 2).
@@
In order for a singularity to develop, the gauge dynamics must drive the disappearance of the IR fixed point, i.e., shift the RG flow parabola such that its zeros merge and vanish. This requires that the leading gauge corrections ($\boldsymbol{c}_{{\rm A},i}$) and the higher-order self-interactions ($\boldsymbol{c}_{{\rm C},iii}$) act coherently, namely by having the same sign, to induce a fixed-point annihilation. While this alignment offers a qualitative criterion for identifying a resonant structure, it relies on a simplified scenario where inter-channel contributions are subleading and the flow is dominated by the leading gauge terms. More intricate dynamics arise when off-diagonal terms become relevant, such as $\boldsymbol{c}_{{\rm B},ij}$ with $i\neq j$ or $\boldsymbol{c}_{{\rm C},ijk}$ with $i\neq j,k$, which can significantly alter the flow and the mechanism of symmetry breaking. These cases will be encountered below.
@@
Moreover, the presence of a resonant structure in a given channel is a necessary but not sufficient condition for a singularity to occur, as the gauge coupling must also reach a sufficient strength to cause the fixed-point merger. To quantify this, we define the critical coupling $\alpha^{\textrm{crit}}_{\textrm{SB}}$ as the minimal value of the gauge coupling $\alpha_{g}$ for which at least one of the four-fermion flows ceases to exhibit a root. Explicitly,
$\alpha^{\textrm{crit}}_{\textrm{SB}}=$ (10)
@@
$\hskip-28.45274pt\inf\left\{\alpha_{g}\,\middle|\,\exists\,i:\,\,\left(\,\partial_{t}\bar{\lambda}_{i}\leq 0\quad\forall\,\bar{\lambda}_{j}\right)\,{\,\rm or\,}\left(\,\partial_{t}\bar{\lambda}_{i}\geq 0\quad\forall\,\bar{\lambda}_{j}\right)\right\}\,,$
@@
which is equivalent to the maximum $\alpha_{g}$ at which all flows remain bounded at a fixed $k$. It follows from (10) that for $\alpha_{g}<\alpha^{\textrm{crit}}_{\textrm{SB}}$, all four-fermion couplings $\bar{\lambda}_{i}$ flow to IR fixed points and for $\alpha_{g}>\alpha^{\textrm{crit}}_{\textrm{SB}}$, at least one coupling ceases to have a fixed point and flows towards a singularity, signaling the onset of dSB. The dominant channel is then identified as the one whose interaction term drives the flow away from the IR fixed point, thereby setting the dominant singularity of the system.
For the class of theories considered here, the diagonal coefficients $\boldsymbol{c}_{{\rm C},iii}$ read:
$\boldsymbol{c}_{{\rm C},\,iii}$ $=\frac{1}{8\pi^{2}}\bigg{\{}\;N_{c}N_{\textrm{gen}}\left(N_{c}-4\right)-1,\;N_{c}+N_{\textrm{gen}}\left(N_{c}-4\right),\;2\,(N_{c}-1)N_{c}+4\,N_{\textrm{gen}},\;\,\frac{(N_{c}-1)N_{c}N_{\textrm{gen}}}{2}-1,$ (11)
$\hskip 213.39566pt\,\frac{N_{c}\,(N_{\textrm{gen}}+3)}{2}+\frac{8}{N_{c}}-\,N_{\textrm{gen}}+1,\;-\frac{3}{2},\;\frac{N_{c}}{4}+\frac{3}{N_{c}}+\frac{3}{4}\bigg{\}}\,.$
@@
Although these coefficients retain a constant sign for $N_{c}>5$, the associated $\boldsymbol{c}_{{\rm A},i}$ undergo a sign change. Recalling the resonance condition, this observation indicates that certain channels may become resonant in specific regions of theory space but not in others, thereby anticipating distinct channel dominance and symmetry-breaking patterns across different domains.
@@
As $N_{c}$ increases, the resonant structure in both ${\cal O}_{5}$ and ${\cal O}_{7}$ is lost due to the sign change in the associated $\boldsymbol{c}_{{\rm A},i}$’s. Around $N_{c}\sim 11$, the simple mechanism ceases to apply in any channel, yet a singularity still develops via a more involved process. For several generations ($N_{\rm gen}\gtrsim 5$ and $N_{c}\gtrsim 11$), off-diagonal contributions $\boldsymbol{c}_{{\rm B},5j}$ and $\boldsymbol{c}_{{\rm C},5jk}$ restore the resonant flow in $\bar{\lambda}_{5}$, effectively leading to the scenario in Region I but with a mild increase in $\alpha^{\textrm{crit}}_{\textrm{SB}}$. The fixed-point merger is now driven by higher-order terms (e.g. $\alpha_{g}\bar{\lambda}_{j}$ and $\bar{\lambda}_{i}\bar{\lambda}_{j}$) rather than the original $g^{4}$. This limit is more non-perturbative, and the key contributions arise from channels such as ${\cal O}_{3}$ and ${\cal O}_{4}$ feeding into ${\cal O}_{5}$. However, as will be discussed below, most of this region likely falls within the conformal window, as the gauge coupling required to trigger symmetry breaking is comparable to or exceeds the values estimated for the IR fixed points.
@@
    Region II: Here, all $\boldsymbol{c}_{{\rm A},i}$ are negative, and the four-fermion couplings flow to large negative values without forming a resonant structure in any single channel. Eventually, all $\bar{\lambda}_{i}\sim\mathcal{O}(1)$, and several couplings diverge simultaneously. In this regime, the system is highly entangled, and no clear dominant channel can be distinguished. We attempted to diagonalize the flow equations to define a basis $\tilde{\lambda}_{i}$ such that each $\tilde{\lambda}_{i}$ is independent of $\tilde{\lambda}_{j}$ for $j\neq i$, but found this not feasible. This regime features a novel mechanism for scale generation driven by high-order interactions, potentially involving exotic higher-dimensional condensates. We note that here the use of the flow equation instead of a perturbative method is particularly beneficial to throw some light into the non-perturbative dSB mechanism. A further alternative scenario in this region concerns the appearance of a gauge-boson mass gap which screens all gauge-fermion interactions. This could prevent the reach of $\alpha^{\textrm{crit}}_{\textrm{SB}}$ and hence the appearance of any condensate, leading to an scenario with color confinement and no global symmetry breaking. The present framework allows for such an assessment Cyrol et al. (2016); Goertz et al. (2024) which will be explored in a future work.
In summary, we found a rich landscape of global symmetry-breaking patterns, which we summarize in Figure 1. The emergence of singularities and resonant channels depends sensitively on $N_{c}$, $N_{\rm gen}$, and the interplay of tensor structures. While bosonization clarifies the nature of the condensates, higher-dimensional operators may be needed to fully capture the dynamics in certain regimes. These results call for future work extending to larger $N_{c}$ and exploring exotic IR phenomena.
@@
By comparing $\alpha^{\textrm{crit}}_{\textrm{SB}}$ with the IR fixed-point values of the gauge beta function, one can estimate the lower boundary of the conformal window, following the approach used for QCD-like systems in Goertz et al. (2024); Gies and Jaeckel (2006). In fact, if the fixed point of the gauge beta function is below the derived $\alpha^{\textrm{crit}}_{\textrm{SB}}$, then the gauge dynamics can never trigger spontaneous symmetry breaking, and eventually flows to an IR-interacting fixed point. In Figure 1, we display the boundaries obtained with the IR fixed point of the two- (dashed), three- (dot-dashed), and four-loop (dotted) $\overline{\text{MS}}$ beta function. The four-loop coefficient, computed for the first time for this class of theories, is presented in Appendix D, along with further details relevant to this regime.
Given the recent interest in these models, particularly for $N_{\textrm{gen}}=1$ Bolognesi et al. (2024, 2020, 2021, 2022); Bolognesi and Konishi (2019); Csáki et al. (2021); Tong (2022); Razamat and Tong (2021); Cacciapaglia et al. (2019), we briefly make contact with the conjectured scenarios. For $N_{c}>5$, the most-attractive-channel arguments Raby et al. (1980); Eichten and Feinberg (1982) and the proposals Bolognesi et al. (2024, 2020, 2021, 2022); Bolognesi and Konishi (2019); Csáki et al. (2021) highlight a first potential $\langle\psi\chi\rangle$ condensate. The latter could break $SU(N_{c})\times SU(N_{c}-4)\rightarrow SU(4)_{c}\times SU(N_{c}-4)_{cf}$ via color-flavor locking, consistent with discrete ’t Hooft anomaly matching conditions. This is in disagreement with our first found condensate $\langle\chi\chi\rangle$, however we identified some clues that a $\langle\psi\chi\rangle$ condensate might also appears at lower energies. Moreover, it is possible that the first $\langle\chi\chi\rangle$ condensate also breaks the gauge group to $SU(4)$, leaving the low-energy theory identical. Finally, contrary to what was found in Csáki et al. (2021), we do not find evidence for the emergence of a $\langle\psi\psi\rangle$ condensate and the respective channels do not show any signature to fulfill the necessary resonant structure.
V Conclusions

### v2

@@ 4 Patterns of symmetry breaking
$\partial_{t}\bar{\lambda}_{i}\propto$  $\,2\,\bar{\lambda}_{i}+\bm{c}_{{\rm A},i}\cdot\alpha_{g}^{2}+\bm{c}_{{\rm B},ij}\cdot\alpha_{g}\,\bar{\lambda}_{j}+\bm{c}_{{\rm C},ijk}\cdot\bar{\lambda}_{j}\bar{\lambda}_{k}+\ldots\,$, (8)
as illustrated in the diagrammatic flow of the four $\psi$-point function sketched in Figure 1. The first term in the right-hand side of (8) corresponds to the canonical scaling of the operator, and the following ones originate from vertex corrections and the fermion anomalous dimension.
  • $\bullet$
    The term $\alpha_{g}^{2}=g^{2}/4\pi$ arises from purely gauge-mediated box diagrams (first diagram on the right-hand side of Figure 1) and is encoded in the coefficient $\bm{c}_{{\rm A},i}$. The coefficients $\bm{c}_{{\rm A},i}$ act as constant shifts and therefore only change the vertical position of the flow without modifying its shape. While the four-fermion interactions are absent at the classical level, these are generated already at one loop by the box diagrams and therefore stress why fermion self-interactions should be interpreted as higher-order gauge corrections, with the various tensor structures corresponding to distinct interaction channels.
  • $\bullet$
    Once induced, these interactions contribute to their own flow through the mixed term $\alpha_{g}\bar{\lambda}_{j}$, which represents a loop involving two gauge vertices and one four-fermion, accounted in $\bm{c}_{{\rm B},ij}$. The coefficients $\bm{c}_{{\rm B},ij}$ contribute as linear corrections act analogously to an anomalous dimension, i.e., as corrections to the canonical scaling.
  • $\bullet$
    Finally, the quadratic term $\bar{\lambda}_{j}\bar{\lambda}_{k}$, encoded in $\bm{c}_{{\rm C},ijk}$, describes purely fermionic self-interactions (right-most diagrams in Figure 1). The indices $j,k$ run over all channels, including $i$, which corresponds to the basis in (6).
The coefficients $\bm{c}_{{\rm A},i}$, $\bm{c}_{{\rm B},ij}$, and $\bm{c}_{{\rm C},ijk}$ encode the color and global symmetries of the theory and exhibit a non-trivial dependence on $N_{c}$ and $N_{\textrm{gen}}$, as determined by traces over the internal structure of the theory.
Moreover, the $\ldots$ in (8) denote further subleading contributions involving one-loop diagrams with other higher dimensional operators. Given the one-loop structure of the flow equation (3), higher $n$-point functions (beyond six) can only contribute indirectly through lower ones. Additional terms with higher powers of $\alpha_{g}$ enter via the anomalous dimensions of the gauge and fermion fields, which appear through the regulator insertions. While these anomalous-dimension effects are included in the present approximation, higher-order corrections originating from six-point functions are not. For further details on this kind of flows, see Braun:2011pp ; Goertz:2024dnz ; Dupuis:2020fhh and references therein.
Figure 2: Depiction of the general form of the flow of a four-fermion coupling, which exhibits a resonant structure as a function of the coupling itself. The red and purple dots indicate an IR-attractive and a UV-repulsive fixed point, respectively. Increasing gauge corrections shift the flow downward, leading to a fixed-point merger at a critical strength $\alpha^{\textrm{crit}}_{\textrm{SB}}$. Beyond this point, the four-fermion coupling runs towards a singularity, signaling the onset of symmetry breaking.
The self-interaction terms $\bm{c}_{{\rm C},iii}$ in (8) shape the RG flow of a given four-fermion coupling as a function of itself into a parabola, see Figures 2 and 6 for an explicit example. Depending on the sign of the self-interaction term (subject to convention), the parabola opens either upwards or downwards. For small $\alpha_{g}$, the flow of the couplings exhibits an IR-attractive fixed point, which in the classical limit ($k\to\infty$) reduces to the Gaussian one, $\alpha_{g}^{*}=0$ and $\bar{\lambda}_{i}^{*}=0$. This IR fixed point [^3] In fact, a partial fixed point of the full system as not all flows vanish (i.e. the gauge coupling flow in theories outside the conformal window) but only those respective to the four-fermion (or higher) interactions. is dictated by the gauge dynamics and reflects once again the interpretation of four-fermion interactions as nothing else but higher-order gauge-fermion interactions. Importantly, the presence of such a fixed point bounds the maximum strength of the fermion self-interactions, thereby preventing the flow from developing a singularity diagnosing dSB.
@@
Furthermore, as the gauge dynamics evolve towards the IR, the parabola shape is shifted upwards or downwards depending on the sign of the $\bm{c}_{{\rm A},i}$ coefficients, which read
$\bm{c}_{{\rm A},\,i}$ $=\frac{9}{2}\bigg\{\;-\frac{1}{4N_{c}^{2}}-\frac{3}{16},\;\,-\frac{N_{c}^{2}-8}{16N_{c}},\;\,-1,\;\,-1+\frac{4+2N_{c}}{N_{c}^{2}},\;\,1-\frac{N_{c}}{8}+\frac{4}{N_{c}},\;\,$
$\hskip 213.39566pt-1+\frac{N_{c}+2}{N_{c}^{2}},\;\,1-\frac{N_{c}}{4}+\frac{4}{N_{c}}\bigg\}\,.$ (9)
@@
Note that none of these depend on $N_{\textrm{gen}}$, but only on $N_{c}$ (see also Figure 4 where we plot these coefficients). For a singularity to develop, the gauge dynamics must drive the disappearance of the IR fixed point, i.e., shift the RG flow parabola such that its zeros merge and vanish into the complex plane. This requires that the leading gauge corrections ($\bm{c}_{{\rm A},i}$) and the higher-order self-interactions ($\bm{c}_{{\rm C},iii}$) act coherently, namely by having the same sign, to induce a fixed-point annihilation. This behavior is characteristic of a resonant structure in the RG flow, cf. (8) and Figure 1, where the respective four-fermion coupling feeds back into its flow constructively. In the absence of such, the lower- and higher-order contributions act incoherently, thereby strengthening the persistence of the fixed point. While the alignment of $\bm{c}_{{\rm A},i}$ and $\bm{c}_{{\rm C},iii}$ having the same sign offers an accurate qualitative criterion for identifying a resonant structure, it assumes a simplified scenario where inter-channel contributions are subleading and the flow is dominated by the leading gauge terms. More intricate dynamics arise when off-diagonal terms become relevant, such as $\bm{c}_{{\rm B},ij}$ with $i\neq j$ or $\bm{c}_{{\rm C},ijk}$ with $i\neq j,k$, which can significantly alter the flow and the mechanism of symmetry breaking. These cases will be encountered below.
@@
Moreover, the presence of a resonant structure in a given channel is a necessary but not sufficient condition for a singularity to occur, as the driving gauge dynamics must also reach a sufficient strength to cause the fixed-point merger. To quantify this, we define the critical coupling $\alpha^{\textrm{crit}}_{\textrm{SB}}$ as the minimal value of $\alpha_{g}$ for which at least one of the four-fermion flows ceases to exhibit an IR fixed-point, namely that $\partial_{t}\bar{\lambda}_{i}<0$ or $\partial_{t}\bar{\lambda}_{i}>0$ uniformly as a function of any $\bar{\lambda}_{j}$. This can be summarized in a compact form in
@@
$\alpha^{\textrm{crit}}_{\textrm{SB}}=\min\left\{\alpha_{g}\,\middle|\,\exists\,i:\,\,\,\left(\,\partial_{t}\bar{\lambda}_{i}\leq 0\quad\forall\,\bar{\lambda}_{j}\right)\,\,{\,\rm or\,}\,\left(\,\partial_{t}\bar{\lambda}_{i}\geq 0\quad\forall\,\bar{\lambda}_{j}\right)\,\,\right\}\,.$ (10)
@@
It then follows, as depicted in Figure 2, that for $\alpha_{g}<\alpha^{\textrm{crit}}_{\textrm{SB}}$, all four-fermion couplings $\bar{\lambda}_{i}$ flow to IR fixed points and for $\alpha_{g}>\alpha^{\textrm{crit}}_{\textrm{SB}}$, at least one coupling ceases to have a fixed point and flows towards a singularity, signaling the onset of dSB. The dominant channel is then identified as the one whose interaction term drives the flow away from the IR fixed point, thereby setting the dominant singularity of the system, that is, a channel whose omission would either eliminate the singularity or raise qualitatively the critical coupling. By contrast, subdominant channels can be removed without affecting the picture. For the class of theories considered here, the diagonal coefficients $\bm{c}_{{\rm C},iii}$ read:
$\bm{c}_{{\rm C},\,iii}$ $=\frac{1}{8\pi^{2}}\bigg\{\;N_{c}N_{\textrm{gen}}\left(N_{c}-4\right)-1,\;N_{c}+N_{\textrm{gen}}\left(N_{c}-4\right),\;2\,(N_{c}-1)N_{c}+4\,N_{\textrm{gen}},\;\,$ (11)
$\hskip 42.67912pt\frac{(N_{c}-1)N_{c}N_{\textrm{gen}}}{2}-1,\,\frac{N_{c}\,(N_{\textrm{gen}}+3)}{2}+\frac{8}{N_{c}}-\,N_{\textrm{gen}}+1,\;-\frac{3}{2},\;\frac{N_{c}}{4}+\frac{3}{N_{c}}+\frac{3}{4}\bigg\}\,.$
@@
Although these coefficients retain a constant sign for $N_{c}>5$, the associated $\bm{c}_{{\rm A},i}$ may undergo a sign change (see Figure 4). According to the resonance condition discussed above, this observation indicates that certain channels may become resonant in specific regions of theory space but not in others, thereby anticipating distinct channel dominance and symmetry-breaking patterns across different domains.
@@
As $N_{c}$ increases, the naive resonant structure in both ${\cal O}_{5}$ and ${\cal O}_{7}$ is lost due to the sign change in the associated $\bm{c}_{{\rm A},i}$’s. Around $N_{c}\sim 11$, the simple mechanism ceases to apply in any channel, yet a singularity still develops via a more involved process. For several generations ($N_{\rm gen}\gtrsim 5$ and $N_{c}\gtrsim 11$), off-diagonal contributions $\bm{c}_{{\rm B},5j}$ and $\bm{c}_{{\rm C},5jk}$ restore the resonant flow in $\bar{\lambda}_{5}$, effectively leading to the scenario in Region I but with a mild increase in $\alpha^{\textrm{crit}}_{\textrm{SB}}$. The fixed-point merger is now driven by higher-order terms (e.g. $\alpha_{g}\bar{\lambda}_{j}$ and $\bar{\lambda}_{i}\bar{\lambda}_{j}$) rather than the original $g^{4}$. This limit is more non-perturbative, and the key contributions arise from channels such as ${\cal O}_{3}$ and ${\cal O}_{4}$ feeding into ${\cal O}_{5}$. However, as will be discussed below, most of this region likely falls within the conformal window, as the gauge coupling required to trigger symmetry breaking is comparable to or exceeds the values estimated for the IR fixed points.
@@
    Region II: Here, all $\bm{c}_{{\rm A},i}$ are negative and therefore none of the channels fulfill the naive condition for the presence of a resonant structure. Nevertheless, we find a singularity. This arises from the interplay between different four-fermion channels in higher-order corrections, rather than being triggered by the primary gauge interactions. As a consequence, the critical coupling $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is shifted to larger values, since higher-dimensional operators must first become large and IR-relevant. This points to a distinct non-perturbative mechanism underlying the initial scale generation which may be very different from the previously discussed and the present in QCD-like theories:
    • Given the large $\alpha^{\textrm{crit}}_{\textrm{SB}}$, it may indicate that multiple bifermion channels turn critical simultaneously and/or that multifermion condensates (encoded in even higher-dimensional operators) are generated.
    • An alternative possibility is that color confinement sets in, generating a mass gap for the gauge modes, before $\alpha^{\textrm{crit}}_{\textrm{SB}}$ is reached. This seems plausible as in QCD-like theories the confining critical coupling is of the same order as the observed here Goertz:2024dnz . In this case, the IR dynamics would be governed by confinement without the formation of fermion condensates, leading to a spectrum of massless baryons needed to match the global anomalies. A specific realization of this idea has been shown to conflict with generalized anomaly matching Bolognesi:2019wfq ; Bolognesi:2020mpe ; Konishi:2024rjz . However, more elaborate realizations, for instance involving distinct sets of massless baryonic states in the IR, could provide a consistent anomaly matching and still remain viable.
    Disentangling such scenarios lies within the reach of functional methods Goertz:2024dnz ; Ferreira:2025anh , and a dedicated study of confinement dynamics in the GG model will be needed to establish whether a similar mechanism occurs there.
In summary, we found a rich landscape of global symmetry-breaking patterns which we summarized in Figure 3. These regions include condensates that break the color group $SU(N_{c})$ similarly to the diquark condensate in the color superconducting phase of QCD Fukushima:2010bq ; Alford:2007xm and which lead to a dynamical Higgs phase with residual gauge dynamics below the first scale. The emergence of divergences and resonant channels depends sensitively on $N_{c}$, $N_{\rm gen}$, and the interplay of tensor structures. Identifying a dominant channel only indicates that a condensate within the fermion content of that channel is expected. More precisely, vacuum expectation values arise in specific fermion combinations organized into irreducible representations. For convenience, we chose a basis $\mathcal{O}_{1,\dots,7}$ in which the flow equations are explicit in $N_{c}$ and $N_{\rm gen}$, which limits a full treatment in terms of irreducible representations. Nevertheless, for particular values of $N_{c}$ and $N_{\rm gen}$ the linear change of basis can be computed, allowing a comparison in the irreducible basis. Even then, however, such an analysis does not fully capture the structure of the vacuum expectation values within the different irreducible condensates. Bosonization, by contrast, can clarify this structure, especially in regimes where higher-dimensional operators may be required to fully describe the dynamics. These results call for future work extending to larger $N_{c}$ and exploring exotic IR phenomena.
@@
By comparing $\alpha^{\textrm{crit}}_{\textrm{SB}}$ with the IR fixed-point values of the gauge beta function, one can estimate the lower boundary of the conformal window, following the approach used for QCD-like systems in Goertz:2024dnz ; Gies:2005as . In fact, if the fixed point of the gauge beta function is below the derived $\alpha^{\textrm{crit}}_{\textrm{SB}}$, then the gauge dynamics can never trigger spontaneous symmetry breaking, and eventually flows to an IR-interacting fixed point. In Figure 3, we display the boundaries obtained with the IR fixed point of the two- (dashed), three- (dot-dashed), and four-loop (dotted) $\overline{\text{MS}}$ beta function. The four-loop coefficient, computed for the first time for this class of theories, is presented in Appendix D, along with further details relevant to this regime.
Last, given the recent interest in these models  Bolognesi:2024bnm ; Bolognesi:2020mpe ; Bolognesi:2021yni ; Bolognesi:2021jzs ; Bolognesi:2019wfq ; Bolognesi:2021hmg ; Csaki:2021xhi ; Tong:2021phe ; Razamat:2020kyf ; Cacciapaglia:2019vce , particularly for $N_{\textrm{gen}}=1$, we now compare the fermion content and symmetry of the condensates. The most-attractive-channel (MAC) approach Raby:1979my ; Eichten:1981mu evaluates the single-gluon exchange between two fermion lines in the irreducible representations and argues for the condensate to appear in the channel with strongest attraction. Our results always disfavor a $\langle\psi\psi\rangle$ condensate, consistent with the MAC. In Region I, we predict a leading $\langle\chi\chi\rangle$ condensate, while the MAC identifies $\langle\chi\chi\rangle$ for $N_{c}=5$, both $\langle\chi\chi\rangle$ and $\langle\psi\chi\rangle$ for $N_{c}=6$, and only $\langle\psi\chi\rangle$ for $N_{c}\geq 7$. Nevertheless, the $\langle\psi\chi\rangle$ channel satisfies the naive resonant condition in part of Region I, suggesting that it may condense at lower energies. Approaches based on anomaly-matching arguments Bolognesi:2024bnm ; Bolognesi:2020mpe ; Bolognesi:2021yni ; Bolognesi:2021jzs ; Bolognesi:2019wfq also highlight the $\langle\psi\chi\rangle$ condensate in a color-flavor-locked vacuum. Remarkably, a $\langle\chi\chi\rangle$ condensate in these scenarios remains consistent with the anomaly conditions. Finally, the analysis from anomaly-mediated-Supersymmetry-breaking Csaki:2021xhi predicts both $\langle\psi\chi\rangle$ and $\langle\psi\psi\rangle$ condensates, where the latter is manifestly disfavored in our findings.
5 Conclusions